3NO2 + 1H2O --> 2HNO3

How many moles of water are needed to react with 8.4 moles of NO2?


1
Expert's answer
2021-01-28T05:36:33-0500

Solution:

The balanced chemical equation:

3NO2 + H2O → 2HNO3 + NO

According to the chemical equation above: n(NO2)/3 = n(H2O)

Hence,

n(H2O) = n(NO2) / 3 = 8.4 mol / 3 = 2.8 mol

n(H2O) = 2.8 mol


Answer: 2.8 moles of water (H2O) are needed to react with 8.4 moles of NO2